OpenStudy (anonymous):

how to factor x^2-10xy+24y^2

OpenStudy (anonymous):

x^2 -10xy + 24y^2 = x^2 -10xy +25y^2 -y^2 = (x - 5y)^2 - y^2 = [(x-5y + y)(x - 5y - y)] = (x - 4y)(x -6y)